OpenStudy (jmartinez638):

How do you find all the solutions for 'sin(x/2 - pi/4) = sqrt2/2'?

OpenStudy (phi):

solutions imply an equation equations have an = sign your expression does not. is that a typo?

OpenStudy (jmartinez638):

Yes, So sorry = sqrt2/2

OpenStudy (jmartinez638):

Edited original.

OpenStudy (phi):

do you know how to "undo" sin ?

OpenStudy (jmartinez638):

Like using identities?

OpenStudy (phi):

more like, do \(\sin^{-1} \) to both sides

OpenStudy (phi):

\[ \sin^{-1} \left(\sin\left(\frac{x}{2} - \frac{\pi}{4}\right) \right)=\sin^{-1} \frac{\sqrt2}{2} \]

OpenStudy (phi):

on the left side, the inverse sin of the sin undoes the sin we are left with \[ \frac{x}{2} - \frac{\pi}{4}=\sin^{-1} \frac{\sqrt2}{2} \]

OpenStudy (phi):

on the right side it is asking for an angle what angle is it where sin of that angle = sqr(2)/2 we want it in radians and it is an angle people memorize (so you should too)

OpenStudy (jmartinez638):

45 degrees or \[\pi\]/4

OpenStudy (phi):

yes. but because the question asks for *all* solutions we should eyeball the graph for sin